### v2

# Connected components in d-minimal structures

Abstract.
For a given d-minimal expansion $\mathfrak{R}$ of the ordered real field, we consider the expansion $\mathfrak{R}^{\natural}$ of $\mathfrak{R}$ generated by the sets of the form $\bigcup_{S\in\mathcal{C}}S$, where $\mathcal{C}$ is a subfamily of the collection of connected components of an $\mathfrak{R}$-definable set. We prove that $\mathfrak{R}^{\natural}$ is d-minimal. A similar assertion holds for almost o-minimal expansions of ordered groups.
Key words and phrases:
d-minimal structures; almost o-minimal structures
2020 Mathematics Subject Classification:
Primary 03C64
The author was supported by JSPS KAKENHI Grant Number JP25K07109. He is grateful to Chris Miller for helpful discussions on this topic.
1. Introduction
Every connected component of a set definable in an o-minimal structure over $\mathbb{R}$ is definable, but this is not the case for locally o-minimal structures and d-minimal structures over $\mathbb{R}$. For instance, [1, Theorem 1.9] gives a necessary and sufficient condition for every connected component of a definable set to be definable for locally o-minimal expansions of (,<,+,)($\mathbb{R}$$\mathbb{Z}$). (,<,+,)($\mathbb{R}$$\mathbb{Z}$) has a definable set whose connected component is not definable. The structure $(\mathbb{R},+,\cdot,2^{\mathbb{Z}})$ is a d-minimal structure by [2], but there exists a set definable in this structure one of whose connected component is not definable [7, Example 79]. A question is whether a given tame structured $\mathfrak{R}$ can be expanded to a structure $\mathfrak{R}^{\natural}$ so that every connected component of a set definable in $\mathfrak{R}$ is definable in $\mathfrak{R}^{\natural}$ without loosing the given tameness such as local o-minimality and d-minimality.
A d-minimal structure [5] is a generalization of o-minimal structures. An expansion $\mathfrak{R}$ of $(\mathbb{R},<)$ is d-minimal if, for every $n\geq 0$ and every definable subset $X$ of $\mathbb{R}^{n+1}$, there exists $N\in\mathbb{N}$ such that the fiber $X_{x}:=\{y\in\mathbb{R}\;|\;(x,y)\in X\}$ is a union of an open set and at most $N$ discrete sets. Our main theorem illustrates how to construct a structure $\mathfrak{R}^{\natural}$ satisfying the above condition when $\mathfrak{R}$ is a d-minimal expansion of an ordered field. The strategy of the proof is applicable to a slightly different setting. An expansion $\mathfrak{R}$ of dense linear order without endpoints is almost o-minimal if every bounded unary definable set is a union of a finite set and finitely many open intervals [3]. As we explain in Remark 3.2, the structure $\mathfrak{R}^{\natural}$ constructed in the same manner is almost o-minimal if $\mathfrak{R}$ is almost o-minimal even if the underlying space is not necessarily $\mathbb{R}$.
For every topological space $T$, $\operatorname{Con}(T)$ denotes the collection of connected components of $T$. For every expansion $\mathfrak{R}$ of $(\mathbb{R},<)$, $\mathfrak{R}^{\natural}$ denotes the expansion of $\mathfrak{R}$ generated by all the sets of the form $\bigcup_{S\in\mathcal{C}}S$ for an $n\in\mathbb{N}$, an $\mathfrak{R}$-definable subset $X$ of $\mathbb{R}^{n}$ and a subfamily $\mathcal{C}$ of $\operatorname{Con}(X)$.
Our main theorem is as follows:
Theorem 1.1.
For every d-minimal expansion $\mathfrak{R}$ of the ordered field of reals, $\mathfrak{R}^{\natural}$ is d-minimal. Furthermore, a subset of $\mathbb{R}^{n}$ is $\mathfrak{R}^{\natural}$-definable if and only if it is a union of finitely many sets of the form $\bigcup_{S\in\mathcal{C}}S$, where $\mathcal{C}$ is a subfamily of $\operatorname{Con}(X)$ for an $\mathfrak{R}$-definable subset $X$ of $\mathbb{R}^{n}$. In particular, for every $\mathfrak{R}^{\natural}$-definable set $X$ and a subfamily $\mathcal{C}$ of $\operatorname{Con}(X)$, $\bigcup_{S\in\mathcal{C}}S$ is $\mathfrak{R}^{\natural}$-definable.
2. Multi-cell decomposition in $\mathfrak{R}$
[…]
The following definition is quoted from [3] and slightly modified. A definable subset $X$ of $\mathbb{R}^{n}$ is a multi-cell if it satisfies the following conditions:
  • If $n=1$, either $X$ is a discrete definable set or all connected components of the definable set $X$ are open intervals.
  • If $n>1$, the projection image $\pi(X)$ is a multi-cell and, for any connected component $Y$ of $X$, $\pi(Y)$ is a connected component of $\pi(X)$ and $Y$ is one of the following forms:
    $Y$ $=\pi(Y)\times\mathbb{R}\text{,}$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y=f(x)\}\text{,}$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y>f(x)\}\text{,}$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;y<g(x)\}\text{ and }$
    $Y$ $=\{(x,y)\in\pi(Y)\times\mathbb{R}\;|\;f(x)<y<g(x)\}$
    for some continuous functions $f$ and $g$ defined on $\pi(Y)$ with $f<g$.
Let $\{A_{1},\ldots,A_{k}\}$ be a nonempty finite family of definable subsets of $\mathbb{R}^{n}$. A multi-cell decomposition partitioning $\{A_{1},\ldots,A_{k}\}$ is a family $\mathcal{C}$ of finitely many multi-cells satisfying the following conditions:
  • For every $1\leq i\leq k$, $A_{i}$ is the union of multi-cells in a subfamily of $\mathcal{C}$;
  • If $n>1$, for every $C_{1},C_{2}\in\mathcal{C}$, either $\pi(C_{1})=\pi(C_{2})$ or $\pi(C_{1})\cap\pi(C_{2})=\emptyset$ holds;
  • If $n>1$, the collection $\pi(\mathcal{C}):=\{\pi(C)\;|\;C\in\mathcal{C}\}$ is a multi-cell decomposition partitioning $\{\pi(A_{1}),\ldots,\pi(A_{k})\}$.
A definable submanifold $M$ of $\mathbb{R}^{n}$ is pseudo-special if $\pi(M)$ is a submanifold of $\mathbb{R}^{n-1}$ and, for every $x\in M$, there exists an open box $B$ in $\mathbb{R}^{n}$ containing $x$ such that $\pi|_{B\cap M}$ is a homeomorphism onto $\pi(B)\cap\pi(M)$. Consider the following property of pseudo-special submanifolds $M$ of $\mathbb{R}^{n}$
  1. ($\dagger$)
    Let $X$ be an arbitrary connected component of $M$. Let $(a_{m})_{m\in\mathbb{N}}$ be a sequence in $\pi(X)$ converging to a point $a\in\pi(M)$. For every $m\in\mathbb{N}$, let $b_{m}\in X$ such that $\pi(b_{m})=a_{m}$. Then, $\{b_{m}\;|\;m\in\mathbb{N}\}$ is bounded.
Observe that every bounded pseudo-special submanifolds enjoys property $(\dagger)$.
Theorem 2.1.
Let $\{A_{1}\ldots,A_{k}\}$ be a nonempty finite family of definable subsets of $\mathbb{R}^{n}$. There exists a multi-cell decomposition partitioning $\{A_{1}\ldots,A_{k}\}$ if one of the following conditions is satisfied:
  1. (a)
    $A_{i}$ is bounded for every $1\leq i\leq k$;
  2. (b)
    For every $n\in\mathbb{N}$ and every pseudo-special submanifold $M$ of $\mathbb{R}^{n}$, property $(\dagger)$ is fulfilled.
The rest of this section is devoted to a proof of Theorem 2.1. This proof is a mixture of Thamrongthanyalak’s proof of [8, Theorem B] and the author’s proof of [3, Theorem 4.22]. Our intermediate target (Lemma 2.4) is to show that a pseudo-special submanifold is a multi-cell if $\pi(M)$ is a multi-cell and $M$ satisfies several additional technical requirements.
We introduce technical definitions. A connected subset $C$ of $\mathbb{R}^{n}$ is a connected graph on a definable subset $V$ of $\mathbb{R}^{n-1}$ if $C$ is the graph of a continuous function $f_{C}:V\to\mathbb{R}$. Throughout, for every connected graph $C$, $f_{C}$ denotes the continuous function whose graph is $C$. For a connected graph $C$ on $V$, we put
$C^{+}:=\{(x,t)\in V\times\mathbb{R}\;|\;t>f_{C}(x)\}\text{ and }$
$C^{-}:=\{(x,t)\in V\times\mathbb{R}\;|\;t<f_{C}(x)\}.$
Let $X$ be a connected subset of $\mathbb{R}^{n}$ with $\pi(X)\subseteq V$ and $X\cap C=\emptyset$. We do not require that $X$ is a connected graph. We have either $X\subseteq C^{+}$ or $X\subseteq C^{-}$. We denote $X>_{V}\underline{C}$ in the former case and $X<_{V}\underline{C}$ in the latter case. Underlines below $C$ are used to clarify which is a connected graph.
Let $C_{1},C_{2}\subseteq\mathbb{R}^{n}$ be two connected graphs on $V$ with $\underline{C_{1}}<_{V}\underline{C_{2}}$. We denote $\{(x,t)\in V\times\mathbb{R}\;|\;f_{C_{1}}(x)<t<f_{C_{2}}(x)\}$ by $(C_{1},C_{2})_{V}$ and $\{(x,t)\in V\times\mathbb{R}\;|\;f_{C_{1}}(x)\leq t\leq f_{C_{2}}(x)\}$ by $C_{1},C_{2}]_{V}$.
Lemma 2.2.
Let $V\subseteq\mathbb{R}^{m-1}$ be a submanifold and $C_{1},C_{2}\subseteq\mathbb{R}^{m}$ be connected graphs on $V$. Let $X$ be a connected subset of $\mathbb{R}^{n}$ such that $\pi(X)$ is a proper open subset of $V$ and $\underline{C_{1}}<_{V}X<_{V}\underline{C_{2}}$. Then $\operatorname{fr}(\pi(X))\cap V\neq\emptyset$, and, for every $x\in\operatorname{fr}(\pi(X))\cap V$, there exists $y\in\operatorname{fr}(X)\cap[C_{1},C_{2}]_{V}$ with $\pi(y)=x$.
Proof.
We may assume that $V$ is bounded by shrinking $V$ if necessary. Connectivity of $V$ and openness of $\pi(X)$ in $V$ imply $\operatorname{fr}(\pi(X))\cap V\neq\emptyset$. Fix an arbitrary $x\in\operatorname{fr}(\pi(X))\cap V$. Replacing $V$ with a small neighborhood of $x$ in $V$, we may assume that $V$ is bounded. Take a sequence $\{x_{m}\}_{m\in\mathbb{N}}$ in $\pi(X)\cap V$ converging to $x$. Let $c_{m}\in\mathbb{R}$ with $(x_{m},c_{m})\in X$. The condition $\underline{C_{1}}<_{V}X<_{V}\underline{C_{2}}$ implies that $X$ is bounded. Since $X$ is bounded, we may assume that $c:=\lim_{m\to\infty}c_{m}$ exists by taking a subsequence. We have $y:=(x,c)\in\operatorname{fr}(X)$. We have $f_{C_{1}}(x)\leq c\leq f_{C_{2}}(x)$ because $\underline{C_{1}}<_{V}X<_{V}\underline{C_{2}}$. This means $y\in[C_{1},C_{2}]_{V}$. ∎
Let $M$ and $P$ be definable subsets of $\mathbb{R}^{n}$. We say that $(M,P)$ is a well-ordered pair if the following conditions are satisfied:
  1. (a)
    $M$ is pseudo-special.
  2. (b)
    $\pi(P)=\pi(M)$;
  3. (c)
    $P\subseteq\operatorname{fr}(M)\cap\pi^{-1}(\pi(M))$;
  4. (d)
    $P$ is a multi-cell.
For a well-ordered pair $(M,P)$, every connected component of the multi-cell $P$ is a connected graph. Let $\operatorname{Adh}^{+}(M,P)$ be the set of points $z\in P$ such that $z\in\operatorname{fr}(M\cap C^{+})$, where $C\in\operatorname{Con}(P)$ with $z\in C$. We define $\operatorname{Adh}^{-}(M,P)$ in the same manner. For every $x\in\pi(M)$, set
$\operatorname{Adh}^{+}(M,P,x):=\{x\}\times\{t\in\mathbb{R}\;|\;(x,t)\in P\cap\operatorname{fr}(M\cap(\{x\}\times(t,\infty)))\}.$
We define $\operatorname{Adh}^{-}(M,P,x)$ similarly. For every $x\in\pi(M)$, we have $\operatorname{Adh}^{+}(M,P,x)\subseteq\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)$ and $\operatorname{Adh}^{-}(M,P,x)\subseteq\operatorname{Adh}^{-}(M,P)\cap\pi^{-1}(x)$.
A well-ordered pair $(M,P)$ is an extremely well-ordered pair if $\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)=\operatorname{Adh}^{+}(M,P,x)$ and $\operatorname{Adh}^{-}(M,P)\cap\pi^{-1}(x)=\operatorname{Adh}^{-}(M,P,x)$ for every $x\in\pi(M)$.
Lemma 2.3.
Let $(M,P)$ be a well-ordered pair. Let $x_{0}\in\pi(M)$. Then, $\operatorname{Adh}^{+}(M,P)$, $\operatorname{Adh}^{-}(M,P)$, $\operatorname{Adh}^{+}(M,P,x_{0})$ and $\operatorname{Adh}^{-}(M,P,x_{0})$ are definable. Furthermore,
$\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P,x)\}$
and
$\{x\in\pi(M)\;|\;\operatorname{Adh}^{-}(M,P)\cap\pi^{-1}(x)\neq\operatorname{Adh}^{-}(M,P,x)\}$
are nowhere dense in $\pi(M)$.
Proof.
We have
$\operatorname{Adh}^{+}(M,P)=\{(x,y)\in\mathbb{R}^{n-1}\times\mathbb{R}\;|\;(x,y)\in P\wedge(\forall B:\text{:open box with }(x,y)\in B$
[…]
$\quad\rightarrow\exists(x_{0},y_{0})\in M\cap B\ (x_{0}\in\pi(P)\cap\pi(B)\wedge y_{0}>f(x_{0}))))\}.$
This implies definability of $\operatorname{Adh}^{+}(M,P)$. $\operatorname{Adh}^{-}(M,P)$ is definable for a similar reason. Definability of $\operatorname{Adh}^{+}(M,P,x_{0})$ and $\operatorname{Adh}^{-}(M,P,x_{0})$ is much easier.
We show that $S:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P)\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P,x)\}$ is nowhere dense in $\pi(M)$. For every $C\in\operatorname{Con}(P)$, let $f_{C}:\pi(C)\to\mathbb{R}$ be the continuous function whose graph is $C$. Let $g_{C}:\pi(C)\to\mathbb{R}$ be the function given by $g_{C}(x):=\inf\{t\in\mathbb{R}\;|\;t>f_{C}(x),\ (x,t)\in M\}$ and
$T_{C}:=\{y\in\operatorname{fr}(M\cap C^{+})\cap P\;|\;f_{C}(\pi(y))<g_{C}(\pi(y))\}.$
Observe that $T_{C}\subseteq C$ and $S=\bigcup_{C\in\operatorname{Con}(P)}\pi(T_{C})$.
Assume for contradiction that $S$ is somewhere dense in $\pi(M)$. Since $S$ is definable in the d-minimal structure $\mathfrak{R}$, $S$ has a nonempty interior in $\pi(M)$ by [5, Main Lemma]. Observe that $P$ has only countably many connected components. By the Baire category theorem, there exists $C\in\operatorname{Con}(P)$ such that $\pi(T_{C})$ is somewhere dense in $\pi(M)$. Take a sufficiently small definable open subset $U$ of $\pi(M)$ such that $\pi^{-1}(U)\cap C$ is definable and $U\cap\pi(T_{C})$ is somewhere dense in $\pi(M)$. This is possible because $P$ is a multi-cell and $\pi(P)=\pi(M)$. Observe that $g_{C}|_{U}$ is definable. Since $U\cap\pi(T_{C})$ is definable in $\mathfrak{R}$, $U\cap\pi(T_{C})$ contains a nonempty definable open subset $V$ of $\pi(M)$ for the same reason as above. We may assume that $g_{C}|_{V}$ is continuous by [5, 8.3 Almost continuity] by shrinking $V$ if necessary. We can easily deduce $\pi(T_{C})\cap V=\emptyset$ because $f_{C}<g_{C}$ on $V$. This is a contradiction. ∎
Let $S$ be a closed subset of a topological space. Let $\operatorname{iso}(S)$ be the set of isolated points in $S$. For every nonnegative integer $m$, we define $\operatorname{iso}_{m}(S)$ as follows:
  • $\operatorname{iso}_{0}(S)=\operatorname{iso}(S)$;
  • $\operatorname{iso}_{m}(S)=\operatorname{iso}(S\setminus\bigcup_{i=0}^{m-1}\operatorname{iso}_{i}(S))$;
Define $\operatorname{rank}(S)$ as the minimum of positive integers $m$ such that $\operatorname{iso}_{m}(S)=\emptyset$. We put $\operatorname{rank}(S)=\infty$ if such an $m$ does not exist.
Lemma 2.4 (Key Lemma).
A pseudo-special submanifold $M$ of $\mathbb{R}^{n}$ satisfying the following conditions is a multi-cell.
  1. (1)
    $M$ enjoys property $(\dagger)$.
  2. (2)
    $\pi(M)$ is a multi-cell.
  3. (3)
    $\operatorname{cl}(M)\cap\pi^{-1}(x)=\operatorname{cl}(M\cap\pi^{-1}(x))$ for every $x\in\pi(M)$;
  4. (4)
    There exists $N\in\mathbb{N}$ with $\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))=N$ for every $x\in\pi(M)$;
  5. (5)
    If $N>1$, for every $1\leq i<N$, the definable set P_i=_x π(M) iso_i(cl(M) π^-1(x)) is a multi-cell. Observe that this condition implies that $(M,P_{i})$ and $(P_{j},P_{k})$ are well-ordered pairs for $1\leq i\leq N-1$ and $1\leq j<k\leq N-1$;
  6. (6)
    $(M,P_{i})$ and $(P_{j},P_{k})$ are extremely well-ordered pairs for $1\leq i\leq N-1$ and $1\leq j<k\leq N-1$.
Proof.
Let $X\in\operatorname{Con}(M)$. Since $X$ is connected, $\pi(X)$ is contained in a connected component of $\pi(M)$, say $V$. Observe that $V$ is simply connected by clause (2) and the definition of multi-cells. We have only to show that $\pi|_{X}$ is a homeomorphism onto $V$ by clause (2). Observe that $\pi|_{X}$ is a local homeomorphism because $\mathcal{M}$ is pseudo-special. First we prove three claims:
Claim 1. Suppose $N=1$. We have $\operatorname{fr}X\cap\pi^{-1}(\pi(M))=\emptyset$.
Proof of Claim 1.
We prove $\operatorname{fr}X\cap\pi^{-1}(x)=\emptyset$ for every $x\in\pi(M)$. The closure of $M\cap\pi^{-1}(x)$ is discrete because $N=1$. This implies that $M\cap\pi^{-1}(x)$ is discrete and closed. We have $\operatorname{fr}X\cap\pi^{-1}(x)\subseteq\operatorname{fr}M\cap\pi^{-1}(x)=\operatorname{fr}(M\cap\pi^{-1}(x))=\emptyset$ by clause (3). ∎
Claim 2. Suppose $N>1$. For every $b\in\operatorname{fr}X\cap\pi^{-1}(\pi(M))$, there exist a definable open subset $V^{\prime}$ of $\pi(M)$ and definable connected graphs $Z^{\prime}$ and $C_{1}$ on $V^{\prime}$ such that $b\in Z^{\prime}$ and one of the following conditions is satisfied:
  • $X>_{V^{\prime}}\underline{Z^{\prime}}$, $\underline{C_{1}}>_{V^{\prime}}\underline{Z^{\prime}}$ and every connected component of $X\cap(Z^{\prime},C_{1})_{V^{\prime}}$ is a connected graph on $V^{\prime}$;
  • $X<_{V^{\prime}}\underline{Z^{\prime}}$, $\underline{C_{1}}<_{V^{\prime}}\underline{Z^{\prime}}$ and every connected component of $X\cap(C_{1},Z^{\prime})_{V^{\prime}}$ is a connected graph on $V^{\prime}$;
In particular, $\pi(b)\in\pi(X)$.
Proof of Claim 2.
Put $a=\pi(b)$. Observe that $b\in\operatorname{fr}X\cap\pi^{-1}(\pi(M))\subseteq\operatorname{fr}M\cap\pi^{-1}(\pi(M))$ by using the assumption that $M$ is pseudo-special. By clauses (3-5),
(1) $\operatorname{fr}M\cap\pi^{-1}(\pi(M))=P_{1}\cup\ldots\cup P_{N-1}$
and $\pi(P_{i})=\pi(M)$ for every $1\leq i<N$. In particular, $V$ is a connected component of $\pi(P_{i})$ for every $1\leq i<N$. There exists $1\leq l<N$ such that $b\in P_{l}$. Set $P:=P_{l}$ and $Q:=\operatorname{fr}M\cap\pi^{-1}(\pi(M))\setminus P$. Let $Z\in\operatorname{Con}(P)$ containing $b$. $Z$ is a connected graph on $V$, and we have either $X>_{V}\underline{Z}$ or $X<_{V}\underline{Z}$. We may assume $X>_{V}\underline{Z}$ without loss of generality. Let $c$ be the last coordinate of $b$. Let $S:=\{s\in\mathbb{R}\;|\;(a,s)\in Q,\ s>c\}$. We consider two separate cases.
Case 1. Suppose $c\in\operatorname{cl}S$. There exists a convergent decreasing sequence $\{s_{m}\}_{m\in\mathbb{N}}$ in $S$ converging to $c$. Since $(a,s_{m})\in Q$, there exists $1\leq k<N$ such that $k\neq l$ and $P_{k}$ contains an infinite subset of $\{(a,s_{m})\;|\;m\in\mathbb{N}\}$. Taking a subsequence, we may assume that $(a,s_{m})\in P_{k}$ for all $m\in\mathbb{N}$. Put $R:=P_{k}$. Let $Y_{m}\in\operatorname{Con}(R)$ containing the point $c_{m}$. Observe that $Y_{m}$ is a connected graph on $V$. We have $f_{Y_{m}}(x)>(c+s_{m})/2$ for every point $x\in V$ sufficiently close to $a$ because $f_{Y_{m}}$ is continuous. Since $b\in\operatorname{fr}X\cap\pi^{-1}(V)$, we can choose $x\in V$ and $t\in\mathbb{R}$ such that $(x,t)\in X$, and $x$ and $t$ are sufficiently close to $a$ and $c$, respectively. This implies that $(x,t)\in Y_{m}^{-}$. We obtain $X<_{V}\underline{Y_{m}}$.
Put $Y^{\prime}:=\operatorname{fr}(\bigcup_{m\in\mathbb{N}}Y_{m})\cap\pi^{-1}(V)$. Observe that $b\in Y^{\prime}$. We show $Y^{\prime}\subseteq Z$. Assume for contradiction that there exists $b^{\prime}\in Y^{\prime}\setminus Z$. Since $b^{\prime}\in\operatorname{fr}M\cap\pi^{-1}(\pi(M))$, there exist $1\leq l^{\prime}<N$ and $R^{\prime}\in\operatorname{Con}(P_{l^{\prime}})$ with $b^{\prime}\in R^{\prime}$ by equality (1). By clause (5), $R^{\prime}$ is a connected graph on $V$. In particular, there exists $c^{\prime}\in\mathbb{R}$ with $(a,c^{\prime})\in R^{\prime}$. Since $X<_{V}\underline{Y_{m}}$ and $\underline{Z}<_{V}X$, we have $\underline{Z}<_{V}\underline{Y_{m}}$ for each $m\in\mathbb{N}$. Since $Z\neq R^{\prime}$, the definition of $b^{\prime}$ and the fact $b^{\prime}\in R^{\prime}$ imply $\underline{R^{\prime}}>_{V}\underline{Z}$. We have $c^{\prime}>c$ and $c^{\prime}<s_{m}$, which contradicts to $c=\lim_{m\to\infty}s_{m}$. We have shown $Y^{\prime}\subseteq Z$.
Pick a point $(x,s)\in X$. Since $X>_{V}\underline{Z}$, there exists $s>u$ with $(x,u)\in Z$. Since $X<_{V}\underline{Y_{m}}$, for every $m\in\mathbb{N}$, there exists $u_{m}>s$ such that $(x,u_{m})\in Y_{m}$. We have $(x,\inf\{u_{m}\;|\;m\in\mathbb{N}\})\in Y^{\prime}\setminus Z$, which is absurd.
Case 2. Next suppose $c\notin\operatorname{cl}(S)$. Since $c\notin\operatorname{cl}(S)$, $c\notin\operatorname{Adh}^{+}(P_{i},P_{l},a)$ for all $1\leq i<l$. We have $b\notin\operatorname{Adh}^{+}(P_{i},P_{l})$ because the pair $(P_{i},P_{l})$ is an extremely well-ordered pair by clause (6). This implies that $b\notin\operatorname{fr}(P_{i}\cap Z_{+})$ for $1\leq i<l$ because $b\in Z$. Consequently, we have $b\notin\bigcup_{i=1}^{N-1}\operatorname{cl}(P_{i}\cap Z^{+})$ by equality (1). On the other hand, we have $\pi^{-1}(V)\cap\operatorname{fr}(M)\cap Z^{+}=\bigcup_{i=1}^{N-1}\pi^{-1}(V)\cap P_{i}\cap Z^{+}\subseteq\bigcup_{i=1}^{N-1}\operatorname{cl}(P_{i}\cap Z^{+})$. Therefore, there exists a bounded open box $B$ containing $b$ with
$\pi(B)\cap\pi(M)\subseteq V$ and $B\cap\operatorname{fr}(M)\cap Z^{+}=\emptyset$.
Let $I$ be an open interval with $B=\pi(B)\times I$.
Since $X>_{V}\underline{Z}$ and $b\in\operatorname{fr}(X)\cap Z$, we have $b\in\operatorname{fr}(M\cap Z^{+})$. This means $b\in\operatorname{Adh}^{+}(M,P)$. We have $b\in\operatorname{Adh}^{+}(M,P,a)$ because $(M,P)$ is an extremely well-ordered pair by clause (6). This means that the closure of $D:=\{t>c\;|\;(a,c)\in M\cap B\}$ contains $c$. Clauses (3,4) imply that $D$ has an empty interior. Take $c_{1}\in D$. Put $c_{m}:=\sup\{t\in D\;|\;t<c_{m-1}\}$ for $m>1$. Since $B\cap\operatorname{fr}(M)\cap Z^{+}=\emptyset$, we have $(a,c_{m})\in M$. Observe that $M\cap(\{a\}\times(c_{m},c_{m-1}))=\emptyset$ for $m>1$. We also have $\lim_{m\to\infty}c_{m}=c$. If not, $\lim_{m\to\infty}c_{m}\in M$ for the same reason as above. This contradicts the assumption that $M$ is pseudo-special. Let $W\in\operatorname{Con}(M)$ with $(a,c_{1})\in W$. Since $M$ is pseudo-special, there exists an open box $B^{\prime}$ in $\mathbb{R}^{n}$ such that $(a,c_{1})\in B^{\prime}$ and $B^{\prime}\cap W$ is the graph of a continuous function defined on $\pi(B^{\prime})\cap\pi(M)$. Shrinking $B^{\prime}$ if necessary, we may assume that $\pi(B^{\prime})\cap\pi(M)=\pi(B^{\prime})\cap V$. Put $V^{\prime}:=\pi(B^{\prime})\cap V$ and $Z^{\prime}=Z\cap\pi^{-1}(V^{\prime})$.
Let $C_{m}\in\operatorname{Con}(M\cap(B^{\prime}\times I))$ containing $(a,c_{m})$ for $m\in\mathbb{N}$. We prove
$\pi(C_{m})=V^{\prime}$
by induction on $m$. Observe that $C_{1}$ is the graph of a continuous function on $V^{\prime}$. Therefore, we have $\pi(C_{1})=V^{\prime}$. Suppose $m>1$. By the induction hypothesis, $C_{m-1}$ is the graph of a continuous function $g_{m-1}$ defined on $V^{\prime}$. Since $c_{m}<c_{m-1}$, $C_{m}$ is connected and $C_{m}\cap C_{m-1}=\emptyset$, we have $C_{m}<_{V^{\prime}}\underline{C_{m-1}}$. Assume for contradiction that $\pi(C_{m})\neq V^{\prime}$. Observe that $V^{\prime}_{m}:=\pi(C_{m})$ is open in $V^{\prime}$ because $M$ is pseudo-special. Since $\underline{Z^{\prime}}<_{V^{\prime}}C_{m}<_{V^{\prime}}\underline{C_{m-1}}$, by Lemma 2.2, $\operatorname{fr}(V^{\prime}_{m})\cap V^{\prime}\neq\emptyset$ and, for every $x^{\prime}\in\operatorname{fr}(V^{\prime}_{m})\cap V^{\prime}$, there exists $y^{\prime}\in\operatorname{fr}(C_{m})\cap[Z^{\prime},C_{m-1}]_{V^{\prime}}$ with $\pi(y^{\prime})=x^{\prime}$. Let $c^{\prime}$ be the last coordinate of $y^{\prime}$. We have $f_{Z}(x^{\prime})\leq c^{\prime}\leq g_{m-1}(x^{\prime})$. In particular, $y^{\prime}=(x^{\prime},c^{\prime})\in B^{\prime}\times I$. We have $y^{\prime}\in\operatorname{fr}(M)$ because $C_{m}\in\operatorname{Con}(M\cap(B^{\prime}\times I))$ and $M$ is pseudo-special. We have $c^{\prime}=f_{Z}(x^{\prime})$ because $B\cap\operatorname{fr}(M)\cap Z^{+}=\emptyset$. In other words, $y^{\prime}\in Z^{\prime}$. Observe that $y^{\prime}\in\operatorname{cl}(M\cap Z^{+})$ because $y^{\prime}\in\operatorname{fr}(C_{m})$ and $\underline{Z^{\prime}}<_{V^{\prime}}C_{m}$. These imply $y^{\prime}\in\operatorname{Adh}^{+}(M,P)$. By clause (5), we have $y^{\prime}\in\operatorname{Adh}^{+}(M,P,x^{\prime})$. This deduces that there exists a point $d^{\prime}\in\mathbb{R}$ with $(x^{\prime},d^{\prime})\in M$ and $c^{\prime}<d^{\prime}<g_{m-1}(x^{\prime})$. Let $C\in\operatorname{Con}(M\cap(B^{\prime}\times I))$ with $(x^{\prime},d^{\prime})\in C$. Since $M$ is pseudo-special, if we choose a sufficiently small open box $B^{\prime\prime}$ containing $(x^{\prime},d^{\prime})$, $M\cap B^{\prime\prime}$ is the graph of a continuous function defined on the submanifold $\pi(M)\cap\pi(B^{\prime\prime})$. We may assume that $\pi(M)\cap\pi(B^{\prime\prime})$ is connected by shrinking $B^{\prime\prime}$ if necessary. Under this condition, we have $M\cap B^{\prime\prime}=C\cap B^{\prime\prime}$ and $\pi(M)\cap\pi(B^{\prime\prime})=\pi(C)\cap\pi(B^{\prime\prime})$. Since $\pi(C)\cap\pi(B^{\prime\prime})$ is a neighborhood of $x^{\prime}$ in $\pi(M)$, $C\cap\pi^{-1}(V^{\prime}_{m})$ is not empty because $x^{\prime}\in\operatorname{fr}(V_{m}^{\prime})$. Pick $C^{\prime}\in\operatorname{Con}(C\cap\pi^{-1}(V^{\prime}_{m}))$. $\pi(C^{\prime})$ is open in $V^{\prime}_{m}$ because $M$ is pseudo-special. We have $\underline{C_{m}}<_{V^{\prime}_{m}}C^{\prime}<_{V^{\prime}_{m}}\underline{(C_{m-1}\cap\pi^{-1}(V^{\prime}_{m}))}$. We also have $\pi(C^{\prime})\neq V^{\prime}_{m}$ because $M\cap(\{a\}\times(c_{m},c_{m-1}))=\emptyset$. By Lemma 2.2, $\operatorname{fr}(C^{\prime})\cap[C_{m},C_{m-1}\cap\pi^{-1}(V^{\prime}_{m})]_{V^{\prime}_{m}}\neq\emptyset$. This implies that $B\cap\operatorname{fr}(M)\cap Z^{+}\neq\emptyset$, which is absurd. We have shown $\pi(C_{m})=V^{\prime}$ for all $m\in\mathbb{N}$.
Next we show
$M\cap(Z^{\prime},C_{1})_{V^{\prime}}=\bigcup_{m>1}C_{m}.$
Let $z\in M\cap(Z^{\prime},C_{1})_{V^{\prime}}\setminus(\bigcup_{m>1}C_{m})$ for contradiction. Put $x_{z}:=\pi(z)$ and $c_{z}$ be the last coordinate of $z$. For $m\in\mathbb{N}$, there exists $d_{m}\in\mathbb{R}$ such that $(x_{z},d_{m})\in C_{m}$ because $\pi(C_{m})=V^{\prime}$. We have $\inf\{d_{m}\;|\;m\in\mathbb{N}\}=f_{Z}(x^{\prime})$ because $B\cap\operatorname{fr}(M)\cap Z^{+}=\emptyset$ and $M$ is pseudo-special. We can choose $m>1$ such that $d_{m}<c_{z}<d_{1}$. Let $C^{\prime\prime}\in\operatorname{Con}(M\cap(Z^{\prime},C_{1})_{V^{\prime}})$ with $z\in C^{\prime\prime}$. We have $\pi(C^{\prime\prime})\neq V^{\prime}$ because $M\cap(\{a\}\times(c,c_{1}))=\{c_{m}\;|\;m>1\}$. By Lemma 2.2, we have $\operatorname{fr}(C^{\prime\prime})\cap[C_{m},C_{1}]_{V^{\prime}}\neq\emptyset$, which contradicts $B\cap\operatorname{fr}(M)\cap Z^{+}=\emptyset$.
Since $M\cap(Z,C_{1})_{V^{\prime}}=\bigcup_{m>1}C_{m}$, every connected component of $X\cap(Z,C_{1})_{V^{\prime}}$ coincides with one of $C_{m}$’s. Therefore, every connected component of $X\cap(Z,C_{1})_{V^{\prime}}$ is a connected graph on $V^{\prime}$. The ‘in particular’ part of Claim 2 is obvious. ∎
Claim 3. Let $b\in X$ and put $a=\pi(b)$.
  1. (a)
    For every continuous map $\alpha:[0,1]\to V^{\prime}$ with $\alpha(0)=a$, there exists a unique continuous map $\beta:[0,1]\to X$ such that $\alpha=\pi\circ\beta$ and $\beta(0)=b$.
  2. (b)
    Let $\beta_{i}:[0,1]\to X$ be continuous maps with $\beta_{i}(0)=a$ for $i=0,1$ and let $H:[0,1]\times[0,1]\to V$ be a homotopy with $H(s,i)=\pi(\beta_{i}(s))$ for $i=0,1$ and $s\in[0,1]$, and $H(0,t)=a$ for $t\in[0,1]$. Then, there exists a unique homotopy $\widetilde{H}:[0,1]\times[0,1]\to X$ between $\beta_{0}$ and $\beta_{1}$ with $\widetilde{H}(0,t)=b$ for $t\in[0,1]$.
Proof of Claim 3.
(a) First we show uniqueness. Let $\beta_{1},\beta_{2}:[0,1]\to X$ be continuous maps with $\alpha=\pi\circ\beta_{i}$ and $\beta_{i}(0)=b$ for $i=1,2$. Since $M$ is a pseudo-special submanifold, $I:=\{t\in[0,1]\;|\;\beta_{1}(t)=\beta_{2}(t)\}$ is a closed and open subset of $[0,1]$ containing 0. We have $I=[0,1]$ because $[0,1]$ is connected.
Consider the set $J$ of $t\in[0,1]$ such that there exists a continuous map $\beta_{t}:[0,t]\to X$ with $\beta_{t}(0)=a$ and $\pi\circ\beta_{t}=\alpha|_{[0,t]}$. $J$ is open because $M$ is pseudo-special. Let $t_{0}:=\sup J$. For every $0\leq t<t_{0}$, there exists a unique continuous map $\beta_{t}:[0,t]\to X$ satisfying the above conditions. The map $\beta:[0,t_{0})\to X$ defined by $\beta(t):=\beta_{t}(t)$ is a continuous map by uniqueness of $\beta_{t}$. Let $(s_{m})_{m\in\mathbb{N}}$ be an increasing sequence in $0,t_{0})$ converging to $t_{0}$. By property $(\dagger)$, we may assume that $(\beta(s_{m}))_{m\in\mathbb{N}}$ is a convergent sequence by taking a subsequence. Put $c=\lim_{m\to\infty}\beta(s_{m})$. We have $\pi(c)=\alpha(t_{0})\in\pi(M)$. Therefore, we obtain $c\in\operatorname{cl}(X)\cap\pi^{-1}(\pi(M))$.
Assume for contradiction that $c\in\operatorname{fr}(X)$. If $N=1$, this contradicts Claim 1. Suppose $N>1$. By Claim 2, there exist a definable open subset $V^{\prime}$ of $\pi(M)$ and definable connected graphs $Z^{\prime}$ and $C_{1}$ on $V^{\prime}$ such that $c\in Z^{\prime}$ and one of the following conditions is satisfied:
  1. (i)
    $X>_{V^{\prime}}\underline{Z^{\prime}}$, $\underline{C_{1}}>_{V^{\prime}}\underline{Z^{\prime}}$ and every connected component of $X\cap(Z^{\prime},C_{1})_{V^{\prime}}$ is a connected graph on $V^{\prime}$;
  2. (ii)
    $X<_{V^{\prime}}\underline{Z^{\prime}}$, $\underline{C_{1}}<_{V^{\prime}}\underline{Z^{\prime}}$ and every connected component of $X\cap(C_{1},Z^{\prime})_{V^{\prime}}$ is a connected graph on $V^{\prime}$;
We may assume that condition (i) is satisfied without loss of generality. By taking $u_{0}<t_{0}$ sufficiently close to $t_{0}$, we may assume that $\alpha([u_{0},t_{0}])$ is contained in $V^{\prime}$. By taking a subsequence of $(s_{m})_{m\in\mathbb{N}}$, we may assume that $\beta(s_{m})\in X\cap(Z^{\prime},C_{1})_{V^{\prime}}$ and $u_{0}<s_{m}<t_{0}$ for every $m\in\mathbb{N}$. By using the assumption that $M$ is a pseudo-special submanifold and condition (i), we can easily prove that $K:=\{t\in[s_{1},t_{0})\;|\;\beta(t)\in X\cap(Z^{\prime},C_{1})_{V^{\prime}}\}$ is closed and open in $s_{1},t_{0})$. We omit the details. Since $s_{1}\in K$, we have $K=[s_{1},t_{0})$. This means that $\beta([s_{1},t_{0}))\subseteq X\cap(Z^{\prime},C_{1})_{V^{\prime}}$. Since $\beta([s_{1},t_{0}))$ is connected, it is contained in a connected component of $X\cap(Z^{\prime},C_{1})_{V^{\prime}}$, say $D$. However, $D$ is a connected graph on $V^{\prime}$ with $Z^{\prime}<_{V^{\prime}}D$ by condition (i). We have $c\in\operatorname{cl}(D)$, but $c\in Z^{\prime}$ and $Z^{\prime}\cap\operatorname{cl}(D^{\prime})=\emptyset$, which is absurd.
We have $c\in X$. This implies $c\in J$. Because $J$ is open in $[0,1]$, we get $c=1$.
(b) Uniqueness follows from clause (a). Clause (a) also implies existence of $\widetilde{H}$ satisfying the requirements in (b) except continuity. Let
$K:=\{s\in[0,1]\;|\;\widetilde{H}|_{[0,s]\times[0,1]}\text{ is continuous}\}.$
Assume for contradiction that $s_{0}:=\sup K<1$. Since $M$ is pseudo-special and uniqueness of $\widetilde{H}$, for every $0\leq t\leq 1$, there exist $\delta_{t}>s_{0}$ and $0\leq\varepsilon_{1,t}<t<\varepsilon_{2,t}\leq 1$ such that the restriction of $\widetilde{H}$ to $0,\delta_{t}]\times[\varepsilon_{1,t},\varepsilon_{2,t}]$ is continuous. Since $[0,1]$ is compact, there exists $\delta>s_{0}$ such that the restriction of $\widetilde{H}$ to $0,\delta]\times[0,1]$ is continuous, which is absurd. ∎
We show $\pi(X)=V$. Since $\pi|_{X}$ is a local homeomorphism, $\pi(X)$ is open in $V$. Assume for contradiction that $\pi(X)$ is not closed in $V$. We can find $a\in\operatorname{fr}(\pi(X))\cap V$. Let $(a_{m})_{m\in\mathbb{N}}$ be a convergent sequence in $\pi(X)$ converging to $a$. Take $b_{m}\in X$ with $\pi(b_{m})=a_{m}$. Since $M$ enjoys property $(\dagger)$, we may assume that $(b_{m})_{m\in\mathbb{N}}$ is convergent in $\mathbb{R}^{n}$ by taking a subsequence. Let $b:=\lim_{m\to\infty}b_{m}$. We have $\pi(b)=a\notin\pi(X)$ and $b\notin X$. This implies that $b\in\operatorname{fr}X\cap\pi^{-1}(\pi(M))$ and $\pi(b)\notin\pi(X)$. These contradict the either Claim 1 or ‘in particular’ part of Claim 2.
The remaining task is to show $\pi|_{X}$ is injective. Let $y_{1},y_{2}\in X$ with $x_{0}:=\pi(y_{1})=\pi(y_{2})$. Let $\gamma:[0,1]\to X$ be a continuous map connecting $y_{1}$ and $y_{2}$. Put $\beta=\pi\circ\gamma$. If $\gamma$ is a constant map, we have $y_{1}=y_{2}$. Therefore, we may assume that $\gamma$ is not a constant map. Observe that $\beta$ is not a constant map because $M$ is pseudo-special. Since $V$ is simply connected, there exists a homotopy $H:[0,1]\times[0,1]\to V$ such that $H(0,t)=\beta(t)$ and $H(1,t)=x_{0}$. By Claim 2, there exists a homotopy $\widetilde{H}:[0,1]\times[0,1]\to X$ such that $\widetilde{H}(0,t)=\gamma(t)$ and $\widetilde{H}(1,t)=y_{1}$. We have $y_{2}=\gamma(1)=\widetilde{H}(1,1)=y_{1}$. ∎
We are now ready to prove Theorem 2.1.
Proof of Theorem 2.1.
We first reduce to the case where $k=1$. Put $A_{i}(0):=A_{i}$ and $A_{i}(1)=\mathbb{R}^{n}\setminus A_{i}$ for $0\leq i\leq k$. Let $\mathcal{S}$ be the set of sequences $\overline{s}$ of zeros and ones of length $k$ with $\overline{s}\neq(1,\ldots,1)$. Set $A_{\overline{s}}:=\bigcap_{i=1}^{k}A_{i}(s_{i})$ for $\overline{s}=(s_{1},\ldots,s_{k})\in\mathcal{S}$. By the hypothesis of reduction, for every $\overline{s}\in\mathcal{S}$, $A_{\overline{s}}$ is partitioned into finitely many multi-cells, say, $C_{\overline{s},1},\ldots,C_{\overline{s},l(\overline{s})}$. If $n=1$, $\{C_{\overline{s},i}\;|\;1\leq i\leq l(\overline{s}),\overline{s}\in\mathcal{S}\}$ is a desired multi-cell decomposition. If $n>1$, there exists a multi-cell decomposition partitioning $\{\pi(C_{\overline{s},i})\;|\;1\leq i\leq l(\overline{s}),\overline{s}\in\mathcal{S}\}$, say $D_{1},\ldots,D_{m}$. Let $\mathcal{D}$ be the collection of nonempty definable sets of the form $C_{\overline{s},i}\cap D_{j}$ for some $\overline{s}\in\mathcal{S}$, $1\leq i\leq l(\overline{s})$ and $1\leq j\leq m$. $\mathcal{D}$ is a desired multi-cell decomposition. Therefore, we may assume that $k=1$. We denote $A_{1}$ by $M$ for simplicity. We use this reduction process in this proof without notice.
We show this theorem by induction on $n$. Suppose $n=1$. Let $J=\operatorname{int}(M)$. Let $N=\operatorname{rank}(\operatorname{cl}(M\setminus J))$ and set $S_{i}:=\operatorname{iso}_{i}(\operatorname{cl}(M\setminus J))\cap(M\setminus J)$ for $0\leq i\leq N-1$. It is easy to show that $\{J,S_{0},\ldots,S_{N-1}\}$ is a multi-cell decomposition of $M$ after removing empty sets in the collection. We omit the details.
Suppose $n>1$. By the induction hypothesis, we can easily reduce to the case where $\pi(M)$ is a multi-cell. This reduction process is also used without notice.
First we consider a special case. Suppose that $\dim M\cap\pi^{-1}(x)=0$ for every $x\in\pi(M)$. We reduce to simpler cases in a step-by-step manner. Set
$N:=\max\{\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))\;|\;x\in\pi(M)\}.$
Set $A^{\prime}:=\bigcup_{x\in\pi(M)}\{x\}\times\operatorname{iso}(\operatorname{cl}(M\cap\pi^{-1}(x)))$. By induction on $N$, by considering $A^{\prime}$ instead of $M$, we may assume that $M\cap\pi^{-1}(x)$ is discrete for every $x\in\pi(M)$.
Put $M[i]:=M\cap\pi^{-1}(\{x\in\pi(M)\;|\;\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))=i\})$ for $1\leq i\leq N$. We may assume that $\operatorname{rank}(\operatorname{cl}(M\cap\pi^{-1}(x)))$ is independent of $x\in\pi(M)$, say $N$, by considering $M[i]$ instead of $M$. We define $P_{k}$ for $1\leq k\leq N-1$ as in clause (5) of Lemma 2.4. Put $P_{0}:=M$ for simplicity of notation. We prove the following claim which implies a multi-cell decomposition partitioning $M$.
Claim 1. There exists a multi-cell decomposition $\{C_{1},\ldots,C_{l}\}$ partitioning $\pi(M)$ such that, for $1\leq i\leq l$, $P_{i}\cap\pi^{-1}(C_{j})$ is a multi-cell for $0\leq i\leq N-1$ and clause (6) in Lemma 2.4 is satisfied for $M\cap\pi^{-1}(C_{i})$.
Proof of Claim 1.
We prove this claim by induction on $N$. First we reduce to a simpler case. There exists a multi-cell decomposition partitioning $\pi(M)$, say $D_{1},\ldots,D_{m}$ by induction hypothesis. By considering $M\cap\pi^{-1}(D_{i})$ instead of $M$, we may assume that $\pi(M)$ is a multi-cell in $\mathbb{R}^{n-1}$ of dimension $d$. Therefore, clause (2) of Lemma 2.4 is always satisfied.
Let $\operatorname{Reg}(M)$ be the set of points $x\in M$ such that there exists an open box $B$ in $\mathbb{R}^{n}$ containing $x$ such that $\pi|_{B\cap M}$ is a homeomorphism onto $\pi(B)\cap\pi(M)$. It is a routine to prove that $\pi(M\setminus\operatorname{Reg}(M))$ has an empty interior in $\pi(M)$. We give a proof here for readers’ convenience. Assume for contradiction, there exists a nonempty definable open subset $U$ of $\pi(M)$ contained in $\pi(M\setminus\operatorname{Reg}(M))$. Since $M\cap\pi^{-1}(x)$ is discrete for $x\in U$, by [6], there exist definable functions $f,g_{1},g_{2}:U\to\mathbb{R}$ such that, for every $x\in U$, $(x,f(x))\in M\setminus\operatorname{Reg}(M)$ and $M\cap(\{x\}\times(g_{1}(x),g_{2}(x)))=\{(x,f(x))\}$. By [5, 8.3 Almost continuity], we may assume that $f,g_{1},g_{2}$ are continuous by shrinking $U$ if necessary. This contradicts that $f(x)\notin\operatorname{Reg}(M)$. Put $S_{1}:=\pi(M\setminus\operatorname{Reg}(M))$, which is nowhere dense in $\pi(M)$.
Put $S_{2}:=\{x\in\pi(M)\;|\;\operatorname{cl}(M)\cap\pi^{-1}(x)\neq\operatorname{cl}(M\cap\pi^{-1}(x))\}$, which is nowhere dense in $\pi(M)$ by [5, Main Lemma(5)]. By induction on $d$, we may assume that $M$ is pseudo-special and clauses (3,4) in Lemma 2.4 are satisfied by considering $M\setminus\pi^{-1}(\operatorname{cl}(S_{1}\cup S_{2}))$ instead of $M$. Observe that clause (1) of Lemma 2.4 always holds by the assumption of this theorem.
First suppose $N=1$. We have $\operatorname{rank}(M\cap\pi^{-1}(x))=1$ for each $x\in\pi(M)$. By the induction hypothesis, there exists a multi-cell decomposition $\{D_{1},\ldots,D_{m}\}$ of $\pi(M)$. We may assume that $\pi(M)$ is a multi-cell by considering $M\cap\pi^{-1}(D_{i})$ instead of $M$. Observe that clauses (5,6) in Lemma 2.4 impose no conditions in this case. $M$ is a multi-cell by Lemma 2.4.
Suppose $N>1$. Apply the induction hypothesis to $P_{1}$, then we may assume that $P_{1},\ldots,P_{N-1}$ are multi-cells and $(P_{i},P_{j})$ are extremely well-ordered pairs for $1\leq i<j\leq N-1$. Observe that $\pi(M)$ is a multi-cell because $\pi(M)=\pi(P_{1})$ and $P_{1}$ is a multi-cell. We want to reduce to the case where $(M,P_{i})$ is a well-ordered pair for every $1\leq i\leq N-1$. Put $T_{1i}:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P_{i})\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P_{i},x)\}$ and $T_{2i}:=\{x\in\pi(M)\;|\;\operatorname{Adh}^{+}(M,P_{i})\cap\pi^{-1}(x)\neq\operatorname{Adh}^{+}(M,P_{i},x)\}$. Put $\bigcup_{i=1}^{N-1}T_{1i}\cup T_{2i}. TT$ is nowhere dense by Lemma 2.3.
We treat this case by induction on $d$. If $d=0$, then $T=\emptyset$. $M$ is a multi-cell by Lemma 2.4. If $d>0$, the claim holds for $M\cap\pi^{-1}(T)$ because $\dim T<d$. Let $\{D_{1},\ldots,D_{l}\}$ be a multi-cell decomposition of $\pi(M)\setminus T$. By considering $M\cap\pi^{-1}(D_{i})$ instead of $M$, we may assume that $T=\emptyset$ and $\pi(M)$ is a multi-cell. We have succeeded in reducing to the case where $M$ is a pseudo-special submanifold enjoying clauses (1) through (6) in Lemma 2.4. Lemma 2.4 implies that $M$ is a multi-cell. ∎
We have proved Theorem 2.1 for the case in which $\dim M\cap\pi^{-1}(x)=0$ for every $x\in\pi(M)$. We treat the general case.
For every set $S\subseteq\mathbb{R}^{n}$, put $S_{x}:=\{y\in\mathbb{R}\;|\;(x,y)\in S\}.$ Consider the set $J=\bigcup_{x\in\pi(M)}\{x\}\times\operatorname{int}(M_{x})$. By Claim 1, there exists a multi-cell decomposition partitioning $M\setminus J$. By considering $J$ instead of $M$, we may assume that $M_{x}$ is open for $x\in\pi(M)$. Consider the following sets:
$V_{+\infty}:=\{(x,y)\in M\;|\;\ \forall t\ (t>y\rightarrow(x,t)\in M)\};$
$V_{-\infty}:=\{(x,y)\in M\;|\;\ \forall t\ (t<y\rightarrow(x,t)\in M)\};$
$V_{\mathbb{R}}:=V_{+\infty}\cap V_{-\infty};$
$W:=M\setminus(V_{+\infty}\cup V_{-\infty}).$
A multi-cell decomposition partitioning $\{V_{+\infty},V_{+\infty},V_{\mathbb{R}},W\}$ is a multi-cell decomposition partitioning $M$. Therefore, we have only to construct multi-cell decompositions partitioning each of $_{+\infty}, V+V_{+\infty}$, $V_{\mathbb{R}}$ and $W$. Observe that $V_{\mathbb{R}}=\pi(V_{\mathbb{R}})\times\mathbb{R}$. It is easy to show that $V_{\mathbb{R}}$ is a union of multi-cells. We omit the details. We next show that there exists a multi-cell decomposition partitioning $V_{+\infty}$. By considering a multi-cell decomposition of $\pi(V_{+\infty})$, we may reduce to the case where $\pi(V_{+\infty})$ is a multi-cell. Consider the definable function $f:\pi(V_{+\infty})\to\mathbb{R}$ given by $f(x)=\inf\{t\in\mathbb{R}\;|\;(x,t)\in V_{+\infty}\}$. Let $D$ be the set of points at which $f$ is discontinuous. By [5, 8.3 Almost continuity], we have $\dim D<\dim\pi(V_{+\infty})$. We can reduce to the case in which $f$ is continuous by induction on $\dim\pi(V_{+\infty})$. $V_{+\infty}$ is a multi-cell because $\pi(V_{+\infty})$ is a multi-cell and $V_{+\infty}=\{(x,t)\in\pi(V_{+\infty})\times\mathbb{R}\;|\;t>f(x)\}$. We can prove that there exists a multi-cell decomposition partitioning $V_{-\infty}$ in the same manner.
The remaining task is to construct a multi-cell decomposition partitioning $W$. Observe that every connected component of $W_{x}$ is a bounded open interval for every $x\in\pi(W)$. As usual, we can reduce to the case in which $\pi(W)$ is a multi-cell by the induction hypothesis. Consider the definable set
$Z:=\{(x,y_{1},y_{2})\in\pi(W)\times\mathbb{R}\times\mathbb{R}\;|\;(x,y_{1})\notin W,(x,y_{2})\notin W,$
$\qquad y_{1}<y_{2},(\forall t\ y_{1}<t<y_{2}\rightarrow(x,t)\in W)\}.$
Let $Q:=\{(x,y)\in\pi(W)\times\mathbb{R}\;|\;\exists z\ (x,y,z)\in Z\}$. Observe that $\pi(W)=\pi(Q)$ and $\dim Q\cap\pi^{-1}(x)=0$ for every $x\in\pi(Q)$. By Claim 1, there exists a multi-cell decomposition $\{Q_{1},\ldots,Q_{l}\}$ of $Q$. Observe that, for every $(x,y)\in Q$, there exists a unique $z\in\mathbb{R}$ with $(x,y,z)\in Z$. Let $g:Q\to\mathbb{R}$ be the definable function sending $(x,y)\in Q$ to the unique point $z\in\mathbb{R}$ with $(x,y,z)\in Z$. Put $g_{i}:=g|_{Q_{i}}$ for $1\leq i\leq l$. By [5, 8.3 Almost continuity], the set $G_{i}$ of points at which $g_{i}$ is discontinuous is nowhere dense in $Q_{i}$. For every $\in\operatorname{Con}(Q_{i}), CC$ is the graph of a continuous function defined on a connected component of $\pi(Q_{i})$. Therefore, $\pi(G_{i}\cap C)$ is nowhere dense in $\pi(Q_{i})$ for every $C\in\operatorname{Con}(Q_{i})$. Since $\operatorname{Con}(Q_{i})$ is a countable set, $\pi(G_{i})=\bigcup_{C\in\operatorname{Con}(Q_{i})}\pi(G_{i}\cap C)$ has an empty interior in $\pi(Q_{i})$ by the Baire category theorem. Since $\pi(G_{i})$ is definable in the d-minimal structure $\mathfrak{R}$, $\dim\pi(G_{i})<\dim\pi(Q_{i})$. By induction on $\dim(\pi(Q_{i}))$, using the induction hypothesis, there exists a multi-cell decomposition $D_{1},\ldots,D_{m}$ of $\pi(Q)$ such that $Q\cap\pi^{-1}(D_{i})$ is a multi-cell and the restriction of $g$ to $Q\cap\pi^{-1}(D_{i})$ is continuous. Since $Q\cap\pi^{-1}(D_{i})$ is a multi-cell, every connected component of $W\cap\pi^{-1}(D_{i})$ is of the form $\{(x,y)\in D_{i}\times\mathbb{R}\;|\;f(x)<y<g(x,f(x))\}$, where $f$ is a continuous function defined on $D_{i}$. This means that $W\cap\pi^{-1}(D_{i})$ is a multi-cell. ∎
3. A generalized version of Theorem 1.1
Theorem 1.1 immediately follows from the following theorem:
Theorem 3.1.
Let $\mathfrak{R}$ be a d-minimal expansion (,<,+,)($\mathbb{R}$$\ldots$) of the ordered group of reals which satisfies one of the following conditions:
  1. (a)
    There exists a definable homeomorphism between $\mathbb{R}$ and a bounded open interval;
  2. (b)
    For every $n\in\mathbb{N}$ and every pseudo-special submanifold $M$ of $\mathbb{R}^{n}$, property $(\dagger)$ is fulfilled.
Every subset $X$ of $\mathbb{R}^{n}$ is $\mathfrak{R}^{\natural}$-definable if and only if there exist a positive integer $l$, $\mathfrak{R}$-definable subsets $Y_{1},\ldots,Y_{l}$ and subfamilies $\mathcal{C}_{i}$ of $\operatorname{Con}(Y_{i})$ such that $X=\bigcup_{i=1}^{l}\bigcup_{S\in\mathcal{C}_{i}}S$.
In particular, $\mathfrak{R}^{\natural}$ is d-minimal and, for every $\mathfrak{R}^{\natural}$-definable set $X$ and a subfamily $\mathcal{C}$ of $\operatorname{Con}(X)$, $\bigcup_{S\in\mathcal{C}}S$ is $\mathfrak{R}^{\natural}$-definable.
Proof.
Let $\mathcal{S}_{n}$ be the collection of subsets $X$ of $\mathbb{R}^{n}$ such that there exist a positive integer $l$, $\mathfrak{R}$-definable subsets $Y_{1},\ldots,Y_{l}$ and subfamilies $\mathcal{C}_{i}$ of $\operatorname{Con}(Y_{i})$ with $X=\bigcup_{i=1}^{l}\bigcup_{S\in\mathcal{C}_{i}}S$. Observe that every $\mathcal{R}$-definable subset of $\mathbb{R}^{n}$ belongs to $\mathcal{S}_{n}$. We may assume that $Y_{1},\ldots Y_{l}$ are multi-cells for the following reason:
We may assume $l=1$. We can decompose $Y_{1}$ into finitely many multi-cells $C_{1},\ldots,C_{k}$ by assumptions (a,b) and Theorem 2.1. For every $1\leq i\leq k$, $D\in\operatorname{Con}(C_{i})$ and $S\in\mathcal{C}_{1}$, we have either $D\subseteq S$ or $D\cap S=\emptyset$ because $D$ and $S$ are connected. Put $\mathcal{D}_{i}=\{D\in\operatorname{Con}(C_{i})\;|\;D\subseteq S\text{ for some }S\in\mathcal{C}_{1}\}$. We have $X=\bigcup_{i=1}^{k}\bigcup_{D\in\mathcal{D}_{i}}D$. By considering $C_{1},\ldots,C_{k}$ and $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$ instead of $Y_{1}$ and $\mathcal{C}_{1}$, we may assume that $Y_{1},\ldots,Y_{l}$ are multi-cells.
It is obvious that every member of $\mathcal{S}_{n}$ is $\mathfrak{R}^{\natural}$-definable. We show that every $\mathfrak{R}^{\natural}$-definable subset of $\mathbb{R}^{n}$ belongs to $\mathcal{S}_{n}$. Since the generators of $\mathfrak{R}^{\natural}$-definable sets belong to $\bigcup_{n=1}^{\infty}\mathcal{S}_{n}$, we have only to prove the following:
  1. (i)
    If $A,B\in\mathcal{S}_{n}$, we have $A\cap B\in\mathcal{S}_{n}$;
  2. (ii)
    If $A\in\mathcal{S}_{n}$, we have $\mathbb{R}^{n}\setminus A\in\mathcal{S}_{n}$.
  3. (iii)
    If $A\in\mathcal{S}_{n}$, we have $\pi(A)\in\mathcal{S}_{n-1}$, where $\pi:\mathbb{R}^{n}\to\mathbb{R}^{n-1}$ denotes the projection forgetting the last coordinate.
By the definition of $\mathcal{S}_{n}$, there exist a positive integer $l_{A}$, $\mathfrak{R}$-definable subsets $Y_{A1},\ldots,Y_{Al_{A}}$ and subfamilies $\mathcal{C}_{Ai}$ of $\operatorname{Con}(Y_{Ai})$ with $A=\bigcup_{i=1}^{l_{A}}\bigcup_{S\in\mathcal{C}_{Ai}}S$. We define $l_{B}$, $Y_{B1},\ldots,Y_{Bl_{B}}$ and $\mathcal{C}_{Bi}$ in the same manner.
We prove clause (i). We can easily reduce to the case where $l_{A}=l_{B}=1$. Let $D$ be an element of $\operatorname{Con}(Y_{A1}\cap Y_{B1})$. For every $S\in\operatorname{Con}(Y_{A1})\cup\operatorname{Con}(Y_{B1})$, we have either $D\subseteq S$ or $D\cap S=\emptyset$ because $D$ and $S$ are connected, $D\subseteq Y_{A1}$ and $D\subseteq Y_{B1}$. Put $\mathcal{C}=\{D\in\operatorname{Con}(Y_{A1}\cap Y_{B1})\;|\;\exists S_{A}\in\mathcal{C}_{A1}\ \exists S_{B}\in\mathcal{C}_{B1}\ D\subseteq S_{A}\cap S_{B}\}$. We have $A\cap B=\bigcup_{D\in\mathcal{C}}D$, which implies $A\cap B\in\mathcal{S}_{n}$.
We prove clause (ii). By clause (i), we may assume $l_{A}=1$. We have $\mathbb{R}^{n}\setminus A=(\mathbb{R}^{n}\setminus Y_{A1})\cup\bigcup_{S\in\operatorname{Con}(Y_{A1})\setminus\mathcal{C}_{A1}}S$. This deduces that $\mathbb{R}^{n}\setminus A\in\mathcal{S}_{n}$.
We prove clause (iii). We may assume $l_{A}=1$ and $Y_{A1}$ is a multi-cell. By the definition of multi-cells, for every connected component of $D$ of $Y_{A1}$, $\pi(D)$ is a connected component of $\pi(Y_{A1})$. Put $\mathcal{E}=\{E\in\operatorname{Con}(\pi(Y_{A1}))\;|\;E=\pi(D)\text{ for some }D\in\mathcal{C}_{A1}\}$. We have $\pi(A)=\bigcup_{E\in\mathcal{E}}E$, which implies $\pi(A)\in\mathcal{S}_{n-1}$.
The remaining task is to show the ‘in particular’ part. We only prove that $\mathfrak{R}^{\natural}$ is d-minimal. The remaining part is easy, and we omit the proof. For every subset $S$ of $\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n-1}$, we denote $\{y\in\mathbb{R}\;|\;(x,y)\in S\}$ by $S_{x}$. Let $X$ be an $\mathcal{R}^{\natural}$-definable subset of $\mathbb{R}^{n}$. We show that there exists a positive integer $m$ such that $X_{x}$ is a union of an open set and at most $m$ discrete sets for every $x\in\mathbb{R}^{n-1}$. There exist a positive integer $l$, $\mathfrak{R}$-definable subsets $Y_{1},\ldots,Y_{l}$ and subfamilies $\mathcal{C}_{i}$ of $\operatorname{Con}(Y_{i})$ with $X=\bigcup_{i=1}^{l}\bigcup_{S\in\mathcal{C}_{i}}S$. We can easily reduce to the case where $l=1$ and, furthermore, we may assume that $Y_{1}$ is a multi-cell by Theorem 2.1. By the definition of multi-cells, $(Y_{1})_{x}$ is either open or discrete for every $x\in\mathbb{R}^{n-1}$. $X_{x}=\bigcup_{S\in\mathcal{C}_{1}}S_{x}$ is also either open or discrete. ∎
Remark 3.2.
An expansion $\mathfrak{R}=(R,<,\ldots)$ of dense linear order without endpoints is called almost o-minimal if every bounded definable subset of $R$ is a union of a finite set and finitely many open intervals. Let $\mathfrak{R}=(R,<,+,\ldots)$ be an almost o-minimal expansion of an ordered group. Note that we do not assume $R=\mathbb{R}$ here and every locally o-minimal structure whose underlying space is $\mathbb{R}$ is almost o-minimal by [3, Proposition 4.11(1)].
We slightly modify the definition of $\mathfrak{R}^{\natural}$ as follows: $\mathfrak{R}^{\natural}$ denotes the expansion of $\mathfrak{R}$ generated by all the sets of the form $\bigcup_{S\in\mathcal{C}}S$ for an $n\in\mathbb{N}$, an $\mathfrak{R}$-definable subset $X$ of $R^{n}$ and a subfamily $\mathcal{C}$ of $\operatorname{Con}_{\text{semi}}(X)$. Here, $\operatorname{Con}_{\text{semi}}(X)$ denotes the collection of semi-definable connected components. Semi-definably connected components of a definable set are defined in [3, Theorem 3.6].
A decomposition theorem into multi-cells holds for almost o-minimal expansion of an ordered group [3, Theorem 4.22]. Using this instead of Theorem 2.1, we can prove the following assertion in the same manner as Theorem 3.1:
A subset $X$ of $R^{n}$ is $\mathfrak{R}^{\natural}$-definable if and only if there exist a positive integer $l$, $\mathfrak{R}$-definable subsets $Y_{1},\ldots,Y_{l}$ and subfamilies $\mathcal{C}_{i}$ of $\operatorname{Con}_{\text{semi}}(Y_{i})$ such that
$X=\bigcup_{i=1}^{l}\bigcup_{S\in\mathcal{C}_{i}}S.$
In particular, for every $\mathfrak{R}^{\natural}$-definable set $X$ and a subfamily $\mathcal{C}$ of $\operatorname{Con}(X)$, $\bigcup_{S\in\mathcal{C}}S$ is $\mathfrak{R}^{\natural}$-definable.
We omit the details of its proof. This assertion implies that $\mathfrak{R}^{\natural}$ is almost o-minimal.